OpenStudy (anonymous):

-2x+4=2(4-3)-3(-8+4x) solve A. X=7 B.x=2 C.x=-7 D.x=3

OpenStudy (anonymous):

-2x+4=8-6+24-12x -2x+12x=8-6+24+4 10x=30 x=30/10 x=3 That is D :)